the lenght of a rectangle is 2 m more than three times its width. If the area is 20 m2 find the diimensions of the rectangle to the nearest hundreth of a meter
Answer by Elomeht(22)

1. The dimensions of the rectangle are x and 3x + 2
2. Since the area is 20, we have x(3x + 2) = 20
3. This becomes (after rearranging) 3x^2 + 2x - 20 = 0
4. Applying the quadratic solution (ignoring negative possibilities), x = (-2 + Square_Root(4 + 240))/6 = 2.27
5. That was the width. The length is 2 + 3 times 2.27 = 2 + 6.81 = 8.81
6. Check: 2.27 times 8.81 = 19.999, which proves that the answers are correct